How many grams of pure iodine are in 2 liters of a 1:20 iodine solution? OA. 10g OB, 40 g ° C. 100 g D. 400 g
Answer #1

answer is c =100 g

pure iodine are in 2 liters of a 1;20 iodine solution

first 2 L change into ml = 2000 ml

let pure iodine present be x

there fore

x/2000 = 1/20

x= 1 x 2000/20

x= 2000/20

x=100

so the answer is 100 g